OpenStudy (anonymous):

Solve the equation. 2x-3y-5z=-16

OpenStudy (anonymous):

Impossible. You need a minimum of 3 equations, or two given variables.

OpenStudy (anonymous):

there's more.. I need to find the n-vector form

OpenStudy (anonymous):

x __ __ __ y = __ + __ s + __ t z __ __ __

OpenStudy (anonymous):

@saifoo.khan can you help me please?

OpenStudy (anonymous):

2x-3y-5z=-16 x+3y+5z-16/2=1 1/2y+2 1/2z-8 that what my calculator said.

OpenStudy (anonymous):

^ huh?

OpenStudy (anonymous):

I put it to my algebra calculator and that the answer I got.

OpenStudy (anonymous):

I'm pretty sure that's wrong ^. But thanks for trying.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

has anyone answered your question?

OpenStudy (anonymous):

and if not can you kindly tell me are you solving for x

OpenStudy (anonymous):

nope, someone tried. I can't figure it out to save my life. In my second comment ( up there ) I'm solving for all the "__"

OpenStudy (anonymous):

ok it is simple

OpenStudy (anonymous):

first of all solve for x then for y then for z

OpenStudy (anonymous):

i will show you how

OpenStudy (anonymous):

you can't solve for x,y,z unless you have more equations..

OpenStudy (anonymous):

wait and you will see

OpenStudy (anonymous):

2x-3y-5z=-16

OpenStudy (anonymous):

step one: 2x = 3y + 5z =16

OpenStudy (anonymous):

step two: \[x = \frac{ 3 }{ 2 }y + \frac{ 5 }{2 }z -8\]

OpenStudy (anonymous):

\[x = 1.5y + 2.5z -8\]

OpenStudy (anonymous):

are you following

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so this is your value of x

OpenStudy (anonymous):

now i will solve for y if you want?

OpenStudy (anonymous):

are you there?